I need these questions answer for these data sets.

(a) Determine the linear prediction rule for predicting empathy (Y) from satisfaction (X). Û = 20.40 + (11.80) (Round to twoPair Number Therapist Empathy Patient Satisfaction 1 67 4 2 89 5 3 41 3 4 44 1 M 60.25 3.25Therapist Empathy and Patient Satisfaction Pair Number Patient Satisfaction Therapist Empathy 72 1 3 2 90 5 3 37 2 4 52 1 M 6Pair Number Patient Satisfaction Therapist Empathy 69 1 4 2 91 5 3 38 2 4 44 1 5 92 1 M 66.8 2.6
Answer #1

y=a+bx b= na (ay) - {xiegi n Ela;)2 – (3x;) a = y bł Ediyi = 792 5601;)? - 39 E ti = 11 n=4 Ey; = 251 ♡ i b= 4(792) - (11) (2

i xi yi (xi)^2 (xi*yi)
1 3 72.0 9 216
2 5 90.0 25 450
3 2 37.0 4 74
4 1 52.0 1 52
Total 11 251.0 39 792.0

100.0 90.0 y = 11.629x + 30.771 R? = 0.7364 80.0 70.0 60.0 50.0 40.0 30.0 20.0 10.0 0.0 1 2 3 4 5 6 X

Si=13 Saiyi - 880 Exi. 51 n=4 EY; = 241 ५ 4 201 4 b- 4(880 - (13)(240) Ib=11.0571 4 (51) - (13) ² 241-(11057) 241 a = 24.3144

i xi yi (xi)^2 (xi*yi)
1 4 67.0 16 268
2 5 89.0 25 445
3 3 41.0 9 123
4 1 44.0 1 44
Total 13 241.0 51 880.0

100.0 90.0 80.0 y = 11.057x + 24.314 R = 0.71 70.0 60.0 50.0 40.0 30.0 20.0 10.0 0.0 1 2 3 4 5 6 X

, Exy = १५० E (1): 47 57013 → 13 ६५ - 33479-334 025 bas(१५3) - (13) (334) →b-5655 5(47) - (15) २ - 3१५ (56515)(१) Q-5211061 -

i xi yi (xi)^2 (xi*yi)
1 4 69.0 16 276
2 5 91.0 25 455
3 2 38.0 4 76
4 1 44.0 1 44
5 1 92.0 1 92
Total 13 334.0 47 943.0

100.0 90.0 80.0 y = 5.6515x + 52.106 R2 = 0.1637 70.0 60.0 > 50.0 40.0 30.0 20.0 10.0 0.0 0 1 2 3 4 5 6 х